OpenStudy (anonymous):

5x^2+x-2+0 quadratic formula

OpenStudy (mathstudent55):

You mean \( 5x^2 + x - 2 = 0\)? If so, use a = 5, b = 1, and c = -2 in the quadratic formula.

OpenStudy (anonymous):

no i meant 3x^2=2x+7

OpenStudy (anonymous):

\[ax^2+bx+c=0 \] is the quadratic formula.

OpenStudy (anonymous):

first i put in other form of.... 4x^2-2x-7=0

OpenStudy (anonymous):

\[3x^2=2x+7\] So subtract 2x and 7... giving you \[3x^2-2x-7=0\]

OpenStudy (anonymous):

hey its 4x^2..sorry 4x^2-2x-7=0

OpenStudy (anonymous):

That's fine. Her is a website that will make all of this much easier for you now and in the future. http://www.mathportal.org/calculators/solving-equations/quadratic-equation-solver.php

OpenStudy (anonymous):

*Here* My bad, typo.

OpenStudy (anonymous):

Okay, sorry. It gives you an explanation. Just though it might help.

OpenStudy (anonymous):

i need someone to walk through with me to see where i go wrong

OpenStudy (anonymous):

@mathstudent55 0

OpenStudy (anonymous):

you have -(-2) but next line as 4.....how?

OpenStudy (mathstudent55):

-(-2) = -1(-2) The product of 2 negative numbers is positive.

OpenStudy (anonymous):

yes but how did it double to 4?

OpenStudy (mathstudent55):

Sorry. My mistake. It's 2. You're right.

OpenStudy (anonymous):

also

OpenStudy (anonymous):

how did 116 go to 29?

OpenStudy (mathstudent55):

I was thinking of the (-2)^2 inside the radical which does turn to 4.

OpenStudy (anonymous):

you can factour the three remaining numbers if necessary and leave in division form

OpenStudy (mathstudent55):

Let me do it agian to correct that mistake and to explain better.

OpenStudy (anonymous):

4x^2=2x+7

OpenStudy (mathstudent55):

\(4x^2 - 2x - 7 = 0\) \(x = \dfrac{-b \pm \sqrt{b^2 - 4ac}} {2a}\) \(x = \dfrac{-(-2) \pm \sqrt{(-2)^2 - 4(4)(-7)}} {2(4)} \) Ok?

OpenStudy (anonymous):

got that as well

OpenStudy (anonymous):

2(+,-) sq rt 4-4(-28) divided by 8

OpenStudy (mathstudent55):

\(x = \dfrac{2 \pm \sqrt{4 - (-112)}} {8}\) \(x = \dfrac{2 \pm \sqrt{4 + 112)}} {8}\) \(x = \dfrac{2 \pm \sqrt{116)}} {8}\) Still good?

OpenStudy (anonymous):

i see, i have the same

OpenStudy (mathstudent55):

Now we need to take care of the 116 indside the radical. \( x = \dfrac{2 \pm \sqrt{116}} {8} \) I have the same.

OpenStudy (mathstudent55):

\( x = \dfrac{2 \pm \sqrt{4 \times 29}} {8} \) \( x = \dfrac{2 \pm \sqrt{4}\sqrt{29}} {8} \) \( x = \dfrac{2 \pm 2\sqrt{29}} {8} \)

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

i dont get that

OpenStudy (anonymous):

how 116 splits up by four and why

OpenStudy (mathstudent55):

116 = 2 * 2 * 29 = 4 * 29 2 is prime and 29 is prime

OpenStudy (anonymous):

ok

OpenStudy (mathstudent55):

The way to simplify a radical is to factor out the largest perfect square you can. The perfect squares that help us here are the squares of the natural numbers: 1, 4, 9, 16, 25, 36, etc.

OpenStudy (anonymous):

give me one to try to split

OpenStudy (mathstudent55):

Since 116 = 4 * 29, 4 is the largest perfect square that can be factored out. The you separate the root into : \(\sqrt{116} = \sqrt{4 \times 29} = \sqrt{4}\sqrt{29} = 2\sqrt{29} \)

OpenStudy (mathstudent55):

I will soon. Let's just finsh this problem now. We're almost done.

OpenStudy (anonymous):

k

OpenStudy (mathstudent55):

We are here: \(x = \dfrac{2 \pm 2\sqrt{29}} {8} \) Now we factor out a 2 from the numerator and a 2 from the denominator.

OpenStudy (anonymous):

1(+,-)sqrt 29 divided by 4

OpenStudy (mathstudent55):

\( x = \dfrac{2(1 \pm \sqrt{29})} {2(4)} \) \( x = \dfrac{2}{2} \times \dfrac{1 \pm \sqrt{29}} {4} \) \( x = \dfrac{1 \pm \sqrt{29}} {4} \) \( x = \dfrac{1 + \sqrt{29}} {4} \) or \( x = \dfrac{1 - \sqrt{29}} {4} \)

OpenStudy (mathstudent55):

Ok?

OpenStudy (anonymous):

yes im stuck on one its6 (+,-)sq rt 24 divided by 6

OpenStudy (mathstudent55):

You mean just to simplify it?

OpenStudy (mathstudent55):

This? \( \dfrac{6 \pm \sqrt{24}}{4} \)

OpenStudy (anonymous):

no all divided by 6

OpenStudy (mathstudent55):

\(\dfrac{6 \pm \sqrt{24}}{6} \) \(=\dfrac{6 \pm \sqrt{4 \times 6}}{6}\)

OpenStudy (anonymous):

is it 3(+.-)sq rt 6 divided by 3?

OpenStudy (mathstudent55):

\(=\dfrac{6 \pm \sqrt{4} \sqrt{6}}{6} \) \(=\dfrac{6 \pm 2 \sqrt{6}}{6} \) \(=\dfrac{3 \pm \sqrt{6}}{3} \) You are correct.

OpenStudy (anonymous):

ok the original problem is 3x^2=6x-1 will you check, sorry

OpenStudy (mathstudent55):

\(3x^2=6x-1\) \(3x^2 - 6x + 1 = 0\)

OpenStudy (mathstudent55):

\(x = \dfrac{-b \pm \sqrt{b^2 - 4ac}} {2a}\) \(x = \dfrac{-(-6) \pm \sqrt{(-6)^2 - 4(3)(1)}} {2(3)}\) Ok?

OpenStudy (mathstudent55):

\(x = \dfrac{6 \pm \sqrt{36 - 12}} {6}\) \(x = \dfrac{6 \pm \sqrt{24}} {6}\) This is what you had, so you are correct.